OpenStudy (deercult):

What are the real and complex solutions of the polynomial equation? 0=x^4+3x^2-4

OpenStudy (marypuro):

so you would input something for x I normally start with 1 and it so happens to work so it would become 0=1^4+3(1)^2-4 1^4=1 3(1)^2=3 1+3=4 4-4=0

OpenStudy (deercult):

Oh, thanks!

OpenStudy (marypuro):

no problem